# Aligning Brain Signals with Multimodal Speech and Vision Embeddings

### v1

@@ 1 Introduction
Human auditory perception is a complex, hierarchical process that begins with the detection of acoustic signals and culminates in the comprehension of spoken language. As sound travels from the ears through auditory pathways, it is incrementally transformed into increasingly abstract representations (phonemes, words, and ultimately meaning), distributed across multiple brain regions [4].
@@
Given this close relationship between auditory stimuli, mental imagery, and neural responses, decoding language from brain signals remains a compelling goal for both neuroscience and artificial intelligence. Prior work by Défossez et al. [3] showed that embeddings from a pre-trained speech model (wav2vec2) could be contrastively aligned with EEG signals, using the average of its final encoder layers.
In this work, we ask a deeper question: which layers of pretrained models best align with neural activity during speech perception? Inspired by the brain’s layered processing, we move beyond averaged representations to perform layer-wise alignment using embeddings from both wav2vec2 and CLIP, the latter offering a lens into the visual associations that may arise when we hear language. We systematically compare three aggregation strategies to evaluate which best aligns with EEG signals recorded while participants listened to a chapter of Alice in Wonderland [2].
@@ 2 Literature Review
Défossez et al. [3] aligned EEG signals with audio representations from a pre-trained wav2vec2 model using a contrastive CLIP-style loss. Their model, built with convolutional and transformer layers, used the average of the final four encoder layers for alignment. While they demonstrated above-chance decoding from EEG, performance was notably higher with MEG data, suggesting limitations in EEG signal quality or modeling.
Our prior work [1] improved EEG decoding by introducing subject-specific attention, personalized spatial mechanisms, and a dual-path RNN, reducing word error rate by up to 1.9%. These results highlighted the importance of architecture design and subject adaptation in brain-to-speech tasks.
@@ 3 Dataset
We use the EEG dataset introduced by Brennan and Hale [2], which includes:
@@ 4 Data Pre-processing
We segmented the EEG recordings into word-level chunks using alignment timestamps provided in the original preprocessing pipeline [2]. For each word, we extended the segment window by 150 milliseconds before and after the onset to account for neural latency.
@@ 5 Methodology
Second, we identify the most predictive layer(s) based on ridge regression performance (using test-set $R^{2}$ and correlation). We then truncate the wav2vec2 model at this optimal layer and use it to generate audio embeddings for training an EEG encoder. This encoder follows the same CNN-transformer architecture as in Défossez et al. [3], and is trained to map EEG data to the selected stimulus embedding space.
@@
For EEG feature extraction, we used the CNN-transformer encoder architecture described in [3]. EEG signals were preprocessed as described earlier and downsampled into fixed-length segments aligned with each word-level audio chunk.

### v2

@@ 1 Introduction
Human auditory perception is a complex, hierarchical process that begins with the detection of acoustic signals and culminates in the comprehension of spoken language. As sound travels from the ears through auditory pathways, it is incrementally transformed into increasingly abstract representations (phonemes, words, and ultimately meaning), distributed across multiple brain regions [holt2022].
@@
Given this close relationship between auditory stimuli, mental imagery, and neural responses, decoding language from brain signals remains a compelling goal for both neuroscience and artificial intelligence. Prior work by Défossez et al. [defossez2023decoding] showed that embeddings from a pre-trained speech model (wav2vec2) could be contrastively aligned with EEG signals, using the average of its final encoder layers.
In this work, we ask a deeper question: which layers of pretrained models best align with neural activity during speech perception? Inspired by the brain’s layered processing, we move beyond averaged representations to perform layer-wise alignment using embeddings from both wav2vec2 and CLIP, the latter offering a lens into the visual associations that may arise when we hear language. We systematically compare three aggregation strategies to evaluate which best aligns with EEG signals recorded while participants listened to a chapter of Alice in Wonderland [brennan_2019].
@@ 2 Literature Review
Défossez et al. [defossez2023decoding] aligned EEG signals with audio representations from a pre-trained wav2vec2 model using a contrastive CLIP-style loss. Their model, built with convolutional and transformer layers, used the average of the final four encoder layers for alignment. While they demonstrated above-chance decoding from EEG, performance was notably higher with MEG data, suggesting limitations in EEG signal quality or modeling.
Our prior work [auster2025penny] improved EEG decoding by introducing subject-specific attention, personalized spatial mechanisms, and a dual-path RNN, reducing word error rate by up to 1.9%. These results highlighted the importance of architecture design and subject adaptation in brain-to-speech tasks.
@@ 3 Dataset
We use the EEG dataset introduced by Brennan and Hale [brennan_2019], which includes:
@@ 4 Data Pre-processing
We segmented the EEG recordings into word-level chunks using alignment timestamps provided in the original preprocessing pipeline [brennan_2019]. For each word, we extended the segment window by 150 milliseconds before and after the onset to account for neural latency.
@@ 5 Methodology
Second, we identify the most predictive layer(s) based on ridge regression performance (using test-set $R^{2}$ and correlation). We then truncate the wav2vec2 model at this optimal layer and use it to generate audio embeddings for training an EEG encoder. This encoder follows the same CNN-transformer architecture as in Défossez et al. [defossez2023decoding], and is trained to map EEG data to the selected stimulus embedding space.
@@
For EEG feature extraction, we used the CNN-transformer encoder architecture described in [defossez2023decoding]. EEG signals were preprocessed as described earlier and downsampled into fixed-length segments aligned with each word-level audio chunk.
@@ Acknowledgment
Annexes
Figure 7: Variance of the First Five PCA Components Across All Stimulus Embeddings
(a) Regression performance (R² and correlation).
(b) Topographic maps of regression weights.
Figure 8: Single-layer regression results (ICA, Subject S04).
(a) Regression performance (R² and correlation).
(b) Topographic maps of regression weights.
Figure 9: Single-layer regression results before feature extraction (PCA, Subject S04).